Your project exceeded costs in the past caused by an underestimation of resource costs in the cost baseline: PV: 51.200.000. EV: 51.000.000. AC: $1.200.000 You expect the underestimation to influence the future as much as it did in the past. If the value of the remaining work (BAC - EV) is at 51.000.000, what should be your new EAC (estimate at completion)? A) $1.800.000 B) $2.000.000 C) $2.200.000 D) $2.400.000
Answer #1

Ans:

Given PV = $ 1,200,000 , EV = $ 1,000,000 , AC = $ 1,200,000

BAC-EV = $ 1,000,000

Need to calculate,

EAC = BAC / CPI

BAC - EV = $ 1,000,000

BAC = 1,000,000 + EV

BAC = $ 2,000,000

CPI = EV / AC

= $ 1,000,000 / $ 1,200,000

= 1 / 1.2

EAC = 2,000,000 * 1.2

= $ 2,400,000

Option D , $ 2,400,000